During one week, Charles swarm for a total of 2.5 h and walked for a total of 8.5 h.

If charles expends 340 kcal/h swimming and 160 kcal/h walking, how many total kilocalories did expend for one week?

please express the energy in kilocalories to two significant figures.
1
Expert's answer
2020-12-22T04:30:33-0500

In a week, Charles;

Swam for 2.5 hours

Walked for 8.5 hours


=> when he swam, the total energy expended = 340kcal/h × 2.5hr = 850kcal

=> when he walked, the total energy expended = 160kcal/h × 8.5hr = 1360kcal


Total energy spent walking and swimming = 1360 + 850 = 2210kcal


This means Charles expended 2200kcal (2 s.f) in one week.